OpenStudy (anonymous):

Line l is perpendicular to the graph of the equation -3x - 4y = 7 and contains the point (4, -4). Find the equation for l.

OpenStudy (anonymous):

-3x-4y=7 rearrange y=-3/4(x)-7/4 m1=-3/4 =original gradient gradient for perpendicular line m1*m2=-1 (-3/4)*m2=-1 m2=4/3 then y=mx+c insert (4,-4) and m2=4/3 -4=(4/3)(4)+c c=-28/3 I equation y=(4/3)x-28/3 or 3y=4x-28

OpenStudy (anonymous):

3y=4x-28 is in the standard form